solve and check
Tickets for a play at the community theater cost $12 for an adult and $8 for a child. If 250 tickets were sold and the total receipts were $2440, how many of each type of ticket were sold?

Answer by ankor@dixie-net.com(12701)

Tickets for a play at the community theater cost $12 for an adult and $8 for a child. If 250 tickets were sold and the total receipts were $2440, how many of each type of ticket were sold?
:
a = number of adult tickets sold
c = number of children's tickets sold
:
a + c = 250
a = (250-c); rearrange for substitution
:
12a + 8c = 2440
Substitute (250-c) for a:
12(250-c) + 8c = 2440
3000 - 12c + 8c = 2440
-4c = 2440 - 3000
-4c = -560
c = -560/-4
c = 140 children
:
250 - 140 = 110 adults
:
Check solution using the $ equation:
12(110) + 8(140) =
1320 + 1140 = 2440